CW-CRDS SPECTRA OF ^{16}O_{3} FROM 5970 TO 6200 CM^{−1}: 2052 TRANSITIONS ASSIGNED AND ANAL-YSED

\underline{A. BARBE}, M-R DE-BACKER-BARILLY, VL. G. TYUTEREV, GSMA, B.P. 1039, 51687 Reims, France; A. CAMPARGUE, S. KASSI, D. ROMANINI, LSP, B.P. 87, 38402 Grenoble, France.

The high resolution CW-CRDS infrared spectra of ozone have been recorded in the 5970-6200 cm^{−1} spectral range. The high sensitivity allow to observe very weak lines. In this region, only two A type bands centred at 6063 and 6124 cm^{−1} were previously analysed^{a}. Here, thanks to the record of new spectra, we largely extend the range of their observed quantum numbers, and, first of all, we observe three B type bands: 2ν_{1} + 3ν_{2} + 2ν_{3} (138 transitions), 5ν_{1} + ν_{2} (16 transitions), and ν_{1} + 2ν_{2} + 4ν_{3} (499 transitions). It is necessary to include 6 vibrational levels in the Hamiltonian model, accounting of many resonances to correctly (r.m.s=6x 10^{−3} cm^{−1} reproduce the observed positions. In addition, transition moment are fitted. Finally, more than 95% of all experimentally observed lines are assigned and are satisfactorily reproduced, accounting of the difficulty of analyses of levels not far from the dissociation.

aM-R De-Backer-Barilly, A. Barbe, Vl G: Tyuterev, D. Romanini, B. Moeskops, A. Campargue, J. Mol. Structure, \underline{780}, 223